### v1

# LitterBox+: An Extensible Framework for LLM-enhanced Scratch Static Code Analysis

Abstract
Large language models (llms) have become an essential tool to support developers using traditional text-based programming languages, but the graphical notation of the block-based Scratch programming environment inhibits the use of llms. To overcome this limitation, we propose the LitterBox+ framework that extends the Scratch static code analysis tool LitterBox with the generative abilities of llms. By converting block-based code to a textual representation suitable for llms, LitterBox+ allows users to query llms about their programs, about quality issues reported by LitterBox, and it allows generating code fixes. Besides offering a programmatic api for these functionalities, LitterBox+ also extends the Scratch user interface to make these functionalities available directly in the environment familiar to learners. The framework is designed to be easily extensible with other prompts, llm providers, and new features combining the program analysis capabilities of LitterBox with the generative features of llms. We provide a screencast demonstrating the tool at https://youtu.be/RZ6E0xgrIgQ.
Index Terms:
Scratch, Block-based Programming, LLM, Automated Feedback
I Introduction
Scratch [1] is a popular block-based programming environment frequently used to introduce beginners to programming [2]. While the block-based nature simplifies coding, beginners often lack a deep understanding of required concepts and introduce issues into their code [3]. llms have recently become popular to help programmers of text-based languages to fix and improve their code, but the graphical nature of block-based code so far hampered the application of llms to Scratch.
To bridge the gap between Scratch programs and the promises of llm support, we introduce the LitterBox+ framework: It uses the LitterBox static code analysis framework [4] to convert Scratch code to a textual representation suitable for llms, and provides a convenient API to query llms about Scratch programs. The framework is designed to be easily extensible with different types of prompts and queries, as well as LLM providers. It automatically provides the context for a given query using the target code in question, offers convenient access to common types of queries, and can build on the existing linting features of LitterBox to allow users to gain further insights and even fix suggestions for the otherwise generic issue descriptions delivered by the static analysis. In addition, LitterBox+ extends the Scratch user interface to enable learners to directly make use of these features (cf. Fig. 1). This includes a custom parser for llm responses such that suggested fixes can be directly applied to Scratch projects.
LitterBox+ is also intended to be an extensible framework upon which researchers can build creative new Scratch+llm integrations, tapping into both the potential of llms as well as the established static analysis features of LitterBox. To support both users and researchers alike, LitterBox+ with its command-line and Java apis as well as the extended Scratch user interface are available as open-source code. The extension to LitterBox is available at https://github.com/se2p/litterbox and our extensions to the Scratch user interface can be found at https://github.com/se2p/NuzzleBug.
Figure 1: A buggy program displayed in the extended Scratch-gui, together with llm controls to explore the bug.
II Background
The different shapes of Scratch blocks determine how they can be combined into programs, thus simplifying programming. However, learners may still struggle to correctly apply programming concepts and produce buggy code. While the process of finding and fixing bugs is an inherent part of learning to program, automated tools can offer support. Since Scratch code is block-based, dedicated analysis tools are required. For example, LitterBox [4] is a linter that automatically detects bugs based on recurring patterns that are common manifestations of misconceptions. Figure 1 shows such a bug pattern in the popular ‘BoatRace’ program [^1] https://scratch.mit.edu/projects/63957956/, accessed 2025-07-21 where a misconception resulted in a ‘Missing Loop’ bug pattern, resulting in a collision detection only being executed once at the start of program execution. Such bug patterns are common in Scratch [3], and feedback on them has been shown to support learners [5]. However, classic static analysis tools like LitterBox can only provide generic help, and learners have to rely on other sources for clarifications or suggestions on how to fix their code.
While llms are capable of providing exactly this help for text-based languages, their application to Scratch code is hampered by the graphical block-based notation. Consequently, existing work on using llms for Scratch tends to focus more on creative aspects, like for example using the llm chat to collect ideas for program features or using the model to generate sprite images [6, 7]. These existing approaches integrate less targeted context, e.g., generic Scratch code patterns from other programs [6], when students ask for support in the llm chat [6, 7]. In contrast, in this paper we focus on providing specific context to the llm in the form of the program’s code, as well as detected issues therein to help the user to solve these issues.
Compared to prior work, we also support more extensive llm-based modifications of existing programs. In particular, we do not rely on heuristic text matching semantics to collect Scratch blocks from the llm’s output [6], but instead developed a parser for the widely used textual scratchblocks [^2] https://en.scratch-wiki.info/wiki/Block_Plugin/Syntax, accessed 2025-07-22 Scratch code representation as part of LitterBox+, which allows us for example to merge fixed scripts returned by the llm into the abstract syntax tree of the surrounding program.
III System Architecture
Figure 2: LitterBox+ architecture overview.
The main interface to our tool for students is via an extension to the Scratch user interface. This direct integration ensures a low barrier to entry by not requiring context switches away from the familiar programming environment. We added a toggleable panel that shows by default issues found by the regular LitterBox static code analysis in the current program (cf. Fig. 1). For each issue, the user can request an extended llm-generated explanation, or request a fix for the issue, which will be automatically integrated into the active program. Additionally, our gui extension offers a chat-like input where a user can ask arbitrary questions about the current sprite or program.
To support the communication between LitterBox running as a Java application on a server and the website in the browser, we extend LitterBox with a rest-api. When we receive a request relating to an issue found by LitterBox from the user interface, we combine information from the issue itself and the program to construct a prompt for the llm. From the issue itself, we extract its type (e.g., Missing Loop) and the default description added by LitterBox. To embed the issue in the relevant context, we also extract the source code of the Scratch program. To provide this code to the llm, we convert the program into the textual scratchblocks format. We use the same format as is used for example on the Scratch community forums and only extend it slightly to include special comments adding unique ids to the start of sprites and scripts. The scratchblocks format was likely seen by the llm during training due to its use on the forums, thus resulting in improved performance over the json program representation. Unlike taking screenshots of the workspace, this removes the visual component from the code. This avoids the problem that code blocks are allowed to overlap in the Scratch coding area; by using the textual format instead of a screenshot of the code, we therefore have more control over which part of the code is included in our prompt.
Depending on the task, we prompt the llm to respond either with an explanation suitable for beginner programmers or respond with a code snippet in scratchblocks format (e.g., when asking for a fix for an issue). To integrate such a code response into the existing program, we add a parser for the scratchblocks format to LitterBox. In case the parsing of a script fails, we extract the script from the response and ask the llm to fix the syntax. If the llm fails to repair the scratchblocks code, the user is informed that the llm returned something unuseful. Since the llm is instructed to retain our special id-comments for modified scripts, we can correlate such code snippets back to the full original program and update the code there accordingly.
Depending on the request, we finally send the response as text, an updated issue, or the program with the llm-proposed fix back to the browser where it can be displayed to the user.
IV User Interfaces
The main user interface is integrated directly into the Scratch web interface. While tests during the implementation of our tool showed that llms can understand and reason about Scratch code, the model may still occasionally produce misleading or wrong explanations and fixes. Since especially learners often cannot discern such cases from helpful ones, we show a prominent hint about the potentially wrong nature in every location where llm-generated content is presented. For example, the red exclamation mark in Fig. 1 shows this hint as tooltip when hovered over. Even though our tool is llm-agnostic, we call all llms GPT in the user interface, since students are likely more familiar with GPT as a general term for all kinds of llms.
By requesting a different output language in the prompts, LitterBox+ can be customised to support any natural language. While the scratchblocks code in the prompts remains English, the Scratch-gui can still automatically display the code in the editor area in the user’s chosen ui language.
For teachers or other more experienced users, we also make all functionality of our LitterBox+ extension available via the command-line interface (cli). Like regular LitterBox, this for example additionally allows for batch-processing of multiple programs or automatically downloading and processing specific programs from the official Scratch website. The extended LitterBox can be started via the cli like a regular Java program:
java -Dlitterbox.llm.openai.api-key=KEY
-jar LitterBox.jar llm --help
For example, to ask the llm to analyse the ‘Boat’ sprite of the program and report new issues missed by LitterBox:
[…].jar llm analyze --path=program.sb3
--target=Boat --new-issues
All subcommands of the cli accept the --help option which explains the available subcommands and their options.
Finally, to allow future research to extend upon our framework, all LitterBox functionality is available via its Java api so that it can also be integrated as a library into other tools. This api allows for future extension in various directions: (1) Our prompt designs can be replaced by customised prompt templates. (2) Future extensions could add support for more model providers besides the currently supported externally hosted (OpenAI) and self-hosted (Ollama) models. [^3] Our implementation uses langchain4j, https://docs.langchain4j.dev/, accessed 2025-07-14. (3) New llm-based features that make use of the existing static code analysers and the scratchblocks parser, or ones based on entirely new analysers could be introduced.
For example, to change the default prompts, only a single ‘prompt provider’ class has to be overridden. It can then be selected as the default prompt provider by setting a system property when starting the LitterBox application. Using a similar selection mechanism, another llm provider can be supported by adding a new concrete implementation of our generic llm-provider interface to LitterBox. The same system properties are used by both LitterBox itself (i.e., when using it as a command-line tool) and by our rest-api extension (i.e., when using it as a backend for the Scratch-gui).
V Features
This section describes the three main features available via the Scratch-gui and other features only available in the cli.
V-A Questions About the Program
Users might have general questions about the behaviour of their program, or they might want to find ideas on how to continue. For this use-case, our framework allows users to ask arbitrary questions about the program in the Scratch user interface. Even though we do not use the static code analysis features in this step, LitterBox is still used to convert the program code into the scratchblocks format. We then combine the user question with this code as relevant context to create the overall prompt for the llm. Depending on the user choice whether the question is specific only to the currently selected sprite or more general about the whole program, only the relevant code context is included in the query to the llm.
V-B Explanation for LitterBox Issues
By default, our extension shows the LitterBox bugs and smells as generated by the static code analysis in the user interface. Their description is mostly generic depending on the bug type, with only some customisation placeholders that are filled for example with the relevant variable names. These generic hints might make it harder for students to come up with reproduction steps or possible fixes since they need to mentally map the generic information to their specific program first. The generic description lacks instructions on how to reproduce the issue, since such steps are always program-specific.
To obtain a more detailed hint, we add a ‘GPT: Explain!’ button to the user interface (cf. Fig. 1). This sends the issue context together with the program to LitterBox. The prompt includes the original issue explanation and asks the llm to (1) explain how the faulty code affects the observable behaviour of sprites and (2) to generate a list of steps required to observe this faulty behaviour.
Since the response is an extended explanation rather than a modified program, it is enough to append the llm’s explanation to the original issue description and send the updated issue back to the browser for display as addition to the previous description (cf. left half of the issue display in Fig. 1). In this example, GPT 4.1444https://openai.com/index/gpt-4-1/, accessed 2025-07-21 correctly explains the issue as follows:
[…]
V-C Automatic Issue Fixing
To extend the support beyond providing an more specific description of the issue, our framework can also propose fixed versions of the code and integrate them directly into the current program. Similar to the explanation, we added a ‘GPT: Fix!’ button into the user interface next to the shown LitterBox issue.
When requesting the llm to propose a suitable fix, we again prompt it with the current issue description and the relevant code context in scratchblocks format. As opposed to the previously described features, the llm now has to produce code in its output rather than only a natural language description. Like for the llm input, we rely on the scratchblocks format and extend LitterBox with a suitable parser. Our parser can process all standard Scratch blocks, custom block definitions, and blocks of the Scratch ‘Pen’ extension. In the future, its grammar could be extended to support further Scratch extensions.
[…]
To integrate the parsed scripts into existing programs, we rely on the special comments marking the beginning of sprites and scripts (cf. Section III). Using the sprite names and block ids contained in these comments, we can align the parsed code snippets with the original code and make the required replacements. These replacements are performed on the internal program representation of LitterBox (i.e., the abstract syntax tree), which is substantially more flexible and reliable compared to working directly on the textual scratchblocks- or json-formats. Any sprites or scripts with unseen ids are added as new code to the program. When adding a new sprite, we use the default image of the Scratch cat mascot.The updated program is finally converted by LitterBox back into the Scratch-native json format and loaded into the Scratch-gui where the user can inspect it in the native code editor (upper half of Fig. 1) and continue to edit it. In case the proposed fix is not satisfactory, the user can choose to revert the change.
V-D Other Features
We also implemented additional features that are currently only available via the cli and the Java api.
V-D1 New Issue Finder
The user can prompt the llm to find additional issues not detected by the LitterBox heuristics in the program. This can also be used to find code ‘perfumes’ [8], i.e., parts of the code following correct programming practices.
V-D2 Code Completion
The user can select a specific script of the program by providing its id and request the llm to extend it with suitable code. The prompt automatically adds the surrounding sprite code as related context to guide the llm to provide useful code that does not duplicate existing parts.
VI Future Work
Since we introduce an initial version of our extensible LitterBox+ framework in this paper, this leaves various opportunities for future work. For use in a classroom setting where the target solution of the program is known, additional agentic capabilities could be integrated that automatically run Whisker [9] or block-based [10] test suites so that either llm-proposed code analysis issue fixes can be checked and improved automatically via a follow-up prompt to the llm, or that fixes for failing tests can be requested directly by the user. Since the Scratch-gui is lacking any kind of code completion feature at the moment, our llm-based completion could be a valuable further extension to the framework and its users. Finally, all features proposed as part of LitterBox+ should be evaluated with children to understand further in how such llm-based tools are used [6, 7] to then adjust prompts and user interface, and develop new features based on the findings.
VII Conclusions
In this paper we introduced our LitterBox+ framework aiming to simplify the integration of llms into Scratch programming tasks. By providing this framework as open-source, we encourage the development of new Scratch+llm integrations going beyond the features presented in this paper.